OpenStudy (anonymous):

Determine which equation represents the following relationship: Ty is 3 years younger than Bea. T = B + 3 T = B - 3 T = 3 - B T = 3B

OpenStudy (anonymous):

@texaschic101 Just a few more :)

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

it is between T=B-3 and T=3-b so it is one of those two

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

it will most likely be T=B-3

OpenStudy (texaschic101):

just write it like you read it... Ty is 3 years younger then Bea.. T = B - 3

OpenStudy (anonymous):

Alright. Thank you! @texaschic101 and @Worior9131

OpenStudy (anonymous):

yo are very welcome